# A Closer Look into Mixture-of-Experts in Large Language Models

### v2

@@ Abstract
Mixture-of-experts (MoE) is gaining increasing attention due to its unique properties and remarkable performance, especially for language tasks. By sparsely activating a subset of parameters for each token, MoE architecture could increase the model size without sacrificing computational efficiency, achieving a better trade-off between performance and training costs. However, the underlying mechanism of MoE still lacks further exploration, and its modularization degree remains questionable. In this paper, we make an initial attempt to understand the inner workings of MoE-based large language models. Concretely, we comprehensively study the parametric and behavioral features of three popular MoE-based models and reveal some intriguing observations, including 1) Neurons act like fine-grained experts; 2) The router of MoE usually selects experts with larger output norms; 3) The expert diversity increases as the layer increases, while the last layer is an outlier, which is further validated by an initial experiment. Based on the observations, we also provide suggestions for a broad spectrum of MoE practitioners, such as router design and expert allocation. We hope this work could shed light on future research on the MoE framework and other modular architectures. Code is available at https://github.com/kamanphoebe/Look-into-MoEs.
@@ A Closer Look into Mixture-of-Experts in Large Language Models
Zili Wang INF Technology                        Jie Fu HKUST
@@ 1 Introduction
Beyond efficiency, another attractive trait of MoE architecture is its modular design and learning paradigm. This modularization allows for flexible and potentially more generalizable handling of diverse data and tasks within a single model by assigning them to specialized experts. Despite its widespread adoption, it remains an open question whether current MoE-based LLMs truly leverage this modularity in knowledge distribution and expert behaviors. In other words, is MoE a simple ensemble of homogeneous experts or a modular combination of heterogeneous experts? Answering this question comprehensively is non-trivial. Therefore, in this paper, we take the first step by investigating three popular MoE-based LLMs (Mixtral 8x7B Jiang et al. (2024), DeepSeekMoE Dai et al. (2024), and Grok-$1$https://github.com/xai-org/grok-1) from two critical perspectives: model parameters and model behaviors. We aim to explore common and distinct features and behaviors among different experts, further shedding light on the inner mechanisms of MoE-based models.
@@
  • Neurons in the Feed-Forward Network (FFN) layer are fine-grained experts. Both the gate embedding and the gate projection matrix of experts perform the choosing operation: the former determines the expert selection while the latter controls the neuron activation. We observe that the similarity heat maps exhibit correlations, suggesting that, from the perspective of the gate projection, the expert neurons can be considered as “tiny” experts, each represented by a single neuron.
@@ 2 Preliminary: Mixture-of-Experts
Mixture-of-Experts models enhance transformers by replacing the original FFNs with $N$ parallel FFNs combined with a router. These $N$ FFNs are also called experts and denoted as $E_{n}$ for $n\in[1,N]$. The router $g(\cdot;\mathbf{G},k)$, parameterized by $\mathbf{G}$ and an integer $\operatorname{k}$, assigns the input $\mathbf{x}$ to a score distribution over the experts, $g(\mathbf{x};\mathbf{G},\operatorname{k})\in\mathbb{R}^{N}$. Typically, the router $g$ consists of a simple linear layer followed by a $\operatorname{softmax}$ and a $\operatorname{Top-k}$ function.
Given $\mathbf{x}\in\mathbb{R}^{h}$, the output $\mathbf{y}\in\mathbb{R}^{h}$ is the weighted sum of the outputs from all experts:
$\mathbf{y}=\sum_{n\in N}g_{n}(\mathbf{x};\mathbf{G},\operatorname{k})E_{n}(\mathbf{x})$
@@
The experts of our chosen models follow the style in LLaMA Touvron et al. (2023), which consists of three linear layers and operates as:
$\text{Expert}(x)=W_{\text{down}}(W_{\text{up}}x\odot\operatorname{Act}(W_{\text{gate}}x))$ (1)
where $\odot$ denotes element-wise multiplication and the activation function is abbreviated as $\operatorname{Act}$. Given the three weight matrices $W_{\text{up}},W_{\text{gate}}\in\mathbb{R}^{d_{\text{mid}}\times d_{\text{hid}}}$ and $W_{\text{down}}\in\mathbb{R}^{d_{\text{hid}}\times d_{\text{mid}}}$, we define a neuron as the combination of the row vectors $W_{\text{up}}[i,:]$ and $W_{\text{gate}}[i,:]$, along with the column vector $W_{\text{down}}[:,i]$. Thus, each expert contains $d_{\text{mid}}$ neurons.
@@ 3 Overview
Model Abbreviation # MoE layers # experts Top-k Hidden size Intermidate size
@@
Our experiments are conducted on several open-source MoE models, namely Mixtral 8x7B, DeepSeekMoE, and Grok-1. We choose these models due to their widespread use and impressive performance across various domains. Additionally, they exhibit complementary characteristics across several key attributes, enabling a robust comparative analysis using control variables. Details are discussed in Append A. To further study the similarities and differences between a standard transformer and a MoE model, we include Mistral 7B Jiang et al. (2023) as one of our investigated models. Basic information about these models, along with the abbreviations used throughout our paper, is summarized in Tab. 1 and Tab. 4. The analysis is divided into two sections: one focusing on model parameters (static) and the other on model behaviors in response to text input (dynamic). Unless otherwise stated, cosine similarity is employed for all experiments involving similarity measurements.
@@ 4.1 Weight Matrices of Experts
MoE models replace FFNs in standard transformers with experts. Following Geva et al. (2020); Qiu et al. (2024b), the projection matrices of the experts can be regarded as keys and values: the column vectors of $W_{\text{down}}$ represent potential outputs; the row vectors of $W_{\text{up}}$ produce weights for each possible output; the row vectors of $W_{\text{gate}}$ determine whether to activate the corresponding neurons. Thus, examining the weight matrices provides a straightforward way to understand the expert behaviors. We analyze both the matrix and neuron levels to gain insights from different perspectives.
@@ 4.1.1 Matrix-level
In this part, we explore the similarity of weight matrices among all experts in each layer for the three projection matrices $W_{\text{up}}$, $W_{\text{gate}}$, and $W_{\text{down}}$. The similarity is calculated based on the flattened matrices and is illustrated in Fig. 1. We denote “F” as the Mistral FFN and “SE” as the DeepSeek shared expert. Note that the figures for different models do not share the same color bar.
@@
DeepSeek and Grok. The shared experts of DeepSeek are implemented as a single MLP block with a larger hidden size than the routed experts, preventing direct comparison of their flattened vectors; thus, we omit them from this experiment. Fig. 1 demonstrates that the similarities between the DeepSeek routed experts and Grok experts are close to zero. While Mixtral’s training method remains unrevealed, it is known that DeepSeek and Grok are trained from scratch. This suggests that Mixtral may have been trained using special schemes, resulting in less diverse experts compared to those trained from scratch Wu et al. (2022).
@@
Figure 1: Matrix-level similarity heat maps of expert weight matrices. Each layer contains three heat maps, corresponding to $W_{\text{up}}$, $W_{\text{gate}}$, and $W_{\text{down}}$, respectively. The tick numbers refer to the expert indices.
@@ 4.1.2 Neuron-level
In § 4.1, we measure the parameter similarity between experts at the matrix level. However, the calculation of cosine similarity is position-dependent. If the neurons of two experts are similar but in different orders, the similarity of their weight matrices will be significantly lower than expected. To address this, we propose two approaches to investigate the correlation at the neuron level: averaging and reordering. Averaging simply averages the rows (for $W_{\text{up}}$ and $W_{\text{gate}}$) or the columns (for $W_{\text{down}}$) of the weight matrices, and then calculates the similarity of the resulting vectors across experts. For reordering, we apply the Jonker-Volgenant algorithm Jonker and Volgenant (1988), which is typically used for solving linear assignment problems, to find the optimal order of neurons that maximizes the cosine similarity between the two experts. We describe the results of the reordering method below and provide the details of the averaging approach in Append D. Additionally, the projection of neurons into low-dimensional spaces using PCA can be found in Append C.2. Due to the heavy computation, we only select several layers for the reordering calculation. Note that the matrices are reordered separately. We measure Kendall’s $\tau$ coefficient between the index sequences before and after reordering, whose value increases when the two sequences exist strong agreement. Tab. 2 depicts the common similarity growth after reordering and the average Kendall’s coefficient $\bar{\tau}$ over the selected layers. The order of Mixtral neurons changes little (resulting in a large $\tau$), and hence nearly unchanged similarities. Despite the substantial similarity increase for DeepSeek and Grok after reordering, their overall values remain around 1e-2.
@@ 4.2 Gate Embedding
The gate in the chosen MoE models is implemented as a linear layer with an embedding size of $\mathbb{R}^{n_{\text{exp}}}\times\mathbb{R}^{d_{\text{hid}}}$, where $n_{\text{exp}}$ is the number of experts. The gate serves as a crucial component of MoE, making it essential to study its attributes to understand MoE functionality better. In addition, since each row vector in the embedding matrix determines expert selection, some correspondence may exist between the embedding matrix and the expert weights.
To investigate this, we measure the similarities between the gate embedding vectors. For computational simplicity, we compare them with the average heat maps (instead of the reordering results) presented in Append D, with qualitative analyses detailed in Append E. Specifically, we found that, for all three MoE models, the patterns in the heat maps of gate vectors and expert neurons in $W_{\text{gate}}$ are partially alike in some layers (i.e., the same coordinates in both heat maps exhibit relatively higher or lower values simultaneously).
@@
Model $(X,Y_{\text{up}})$ $(X,Y_{\text{gate}})$ $(X,Y_{\text{down}})$
@@ 4.3 Summary
Here, we conclude the key observations from the analysis of static parameters: i) Mixtral might contain expert(s) with unique attributes, as evidenced by the frequent presence of dark crosses in Fig. 1. ii) The similarities of DeepSeek and Grok expert weight matrices are generally lower than those in Mixtral. As mentioned in § 4.1.1, the matrix-level similarities of DeepSeek and Grok experts are typically close to zero, whereas Mixtral expert similarities average around 0.3. iii) The weights of different experts become less similar in deeper layers, as observed in the Mixtral heat maps in Fig. 1. iv) $W_{\text{up}}$, $W_{\text{down}}$, and $W_{\text{gate}}$, share similar patterns in their similarity heat maps (Fig. 1). v) The similarities of the gate embeddings and of $W_{\text{gate}}$ show either positive or negative association. Tab. 3 depicts the $R_{\text{avg}}^{2}$ values, where the pairing of gate embeddings and $W_{\text{gate}}$ achieves the highest correlation across all three models.
@@ 5 Analysis of Dynamic Behaviours
Refer to caption      Refer to caption
Refer to caption      Refer to caption
Refer to caption Refer to caption
Refer to caption     Refer to caption
Figure 2: Similarity heat maps of expert output features using the short input. The top $k$ experts for each token are shown on top of each heat map. The tick numbers refer to the expert indices.
@@
Figure 3: Average similarity heat maps of expert output features using the long input, plotted along with the matrix-level similarity heat maps. The tick numbers refer to the expert indices.
@@
We utilize a short input and a long input for the experiments in this section. For the short input, we employ the first few words of the input from another MoE-related work Cai et al. (2024)  [^3] The specific tokens are ¡s¿, As, an, open, source, alternative, to, where the start of the sentence symbol ¡s¿ does not applicable for the Grok tokenizer.. For the long input, we adopt 10 sequences from the test set of the WikiText-103 Merity et al. (2016) dataset, totaling approximately 1100 tokens. The sequences in WikiText-103 cover a variety of domains, with the 10 sequences we used spanning topics such as music, design, and construction. To ensure the robustness of our findings, we repeat experiments requiring the long input (§ 5.1, § 5.2) using additional datasets with over 80K tokens. Observations align with the results described in the main context, even with subject-specific datasets like math and code. See Append F for dataset details.
@@ 5.1 Outputs of Experts
For clarity, the average similarity heat maps are plotted alongside the neuron-level similarity graphs of the expert weight matrices. Fig. 9 further depicts the results from additional datasets, which are consistent with those of the long input.
@@
DeepSeek. Given the short input, most similarities are around zero, while the values in the last layer are significantly larger. Again, the similarities between experts chosen by the gate are likely to be higher, although this difference occurs much less frequently than in Mixtral. The average similarities for the long input also approach zero. Moreover, the number of “small rectangular” with relatively light color in the graphs decreases as the layer depth increases (except for the last layer), meaning that the average similarities gradually decline.
Grok. Surprisingly, the similarities between the output features remain high for all tokens in the short input, indicating the experts exhibit similar behaviours. However, the similarities of their weight matrices are mostly zeros (§ 4.1.1). We speculate that this may be due to the relatively large size of each Grok expert, allowing each to learn comprehensive knowledge and behave similarly despite having distinct parameters. When averaging the similarities for the long input, some of the resulting average heat maps display patterns similar to those of the $W_{\text{gate}}$ figures. This relationship aligns with the observations made for Mixtral.

### v3

@@ Abstract
Mixture-of-experts (MoE) is gaining increasing attention due to its unique properties and remarkable performance, especially for language tasks. By sparsely activating a subset of parameters for each token, MoE architecture could increase the model size without sacrificing computational efficiency, achieving a better trade-off between performance and training costs. However, the underlying mechanism of MoE still lacks further exploration, and its modularization degree remains questionable. In this paper, we make an initial attempt to understand the inner workings of MoE-based large language models. Concretely, we comprehensively study the parametric and behavioral features of four popular MoE-based models and reveal some intriguing observations, including 1) Neurons act like fine-grained experts; 2) The router of MoE usually selects experts with larger output norms; 3) The expert diversity increases as the layer increases, while the last layer is an outlier, which is further validated by an initial experiment. Based on the observations, we also provide suggestions for a broad spectrum of MoE practitioners, such as router design and expert allocation. We hope this work could shed light on future research on the MoE framework and other modular architectures. Code is available at https://github.com/kamanphoebe/Look-into-MoEs.
@@ A Closer Look into Mixture-of-Experts in Large Language Models
Zili Wang INF Technology                        Jie Fu Shanghai AI Lab
@@ 1 Introduction
Beyond efficiency, another attractive trait of MoE architecture is its modular design and learning paradigm. This modularization allows for flexible and potentially more generalizable handling of diverse data and tasks within a single model by assigning them to specialized experts. Despite its widespread adoption, it remains an open question whether current MoE-based LLMs truly leverage this modularity in knowledge distribution and expert behaviors. In other words, is MoE a simple ensemble of homogeneous experts or a modular combination of heterogeneous experts? Answering this question comprehensively is non-trivial. Therefore, in this paper, we take the first step by investigating four popular MoE-based LLMs (Mixtral 8x7B Jiang et al. (2024), Mixtral 8x22B, DeepSeekMoE Dai et al. (2024), and Grok-$1$https://github.com/xai-org/grok-1) from two critical perspectives: model parameters and model behaviors. We aim to explore common and distinct features and behaviors among different experts, further shedding light on the inner mechanisms of MoE-based models.
@@
  • Neurons in the Feed-Forward Network (FFN) layer are fine-grained experts. Both the gate embedding matrix and the expert projection matrix $W_{\text{act}}$ perform the choosing operation: the former determines the expert selection while the latter controls the neuron activation. We observe that the similarity heat maps exhibit correlations, suggesting that, from the perspective of $W_{\text{act}}$, the expert neurons can be considered as “tiny” experts, each represented by a single neuron.
@@ 2 Preliminary: Mixture-of-Experts
Mixture-of-Experts models enhance transformers by replacing the original FFNs with $N$ parallel FFNs combined with a router. These $N$ FFNs are called experts and denoted as $E_{n}$ for $n\in[1,N]$. The gate $g(\cdot;W_{g},k)$, parameterized by $W_{g}$ and an integer $\operatorname{k}$, assigns the input $x$ to a score distribution over the experts, $g(x;W_{g},\operatorname{k})\in\mathbb{R}^{N}$. Typically, the gate $g$ consists of a simple linear layer followed by a $\operatorname{softmax}$ and a $\operatorname{Top-k}$ function.
Given $x\in\mathbb{R}^{d_{\text{hid}}}$, the output $y\in\mathbb{R}^{d_{\text{hid}}}$ is the weighted sum of the outputs from all experts:
$y=\sum_{n\in N}g_{n}(x;W_{g},\operatorname{k})E_{n}(x)$
@@
The experts $E_{n}$ of the models investigated in this paper follow the style in LLaMA Touvron et al. (2023), which consists of three linear layers and operates as (the subscript $n$ is omitted for brevity):
$E(x)=W_{\text{down}}(W_{\text{up}}x\odot\sigma(W_{\text{act}}x))$ (1)
where $\odot$ represents element-wise multiplication and $\sigma$ represents the activation function. Given the three projection matrices $W_{\text{up}},W_{\text{act}}\in\mathbb{R}^{d_{\text{mid}}\times d_{\text{hid}}}$ and $W_{\text{down}}\in\mathbb{R}^{d_{\text{hid}}\times d_{\text{mid}}}$, we define a neuron as the combination of the row vectors $W_{\text{up}}[i,:]$ and $W_{\text{act}}[i,:]$, along with the column vector $W_{\text{down}}[:,i]$. Thus, each expert contains $d_{\text{mid}}$ neurons, each with size $d_{\text{hid}}$.
@@ 3 Overview
Model Abbreviation # MoE layers # experts Top-k Hidden size Intermediate size
@@
Our experiments are conducted on several open-source MoE models, namely Mixtral 8x7B, Mixtral 8x22B  [^2] For the Mixtral 8x22B model, we only conduct most of the analyses mentioned in the main context (excluding those in the appendix) due to time limit., DeepSeekMoE, and Grok-1. We choose these models due to their widespread use and impressive performance across various domains. Additionally, they exhibit complementary characteristics across several key attributes, enabling a robust comparative analysis using control variables. Details are discussed in Append A. To further study the similarities and differences between a standard transformer and a MoE model, we include Mistral 7B Jiang et al. (2023) as one of our investigated models. Basic information about these models, along with the abbreviations used throughout our paper, is summarized in Tab. 1 and Tab. 4. The analysis is divided into two sections: one focusing on model parameters (static) and the other on model behaviors in response to text input (dynamic).
Unless otherwise stated (§ 5.1), cosine similarity is employed for all experiments involving similarity measurements. While we acknowledge the existence of other metrics, we primarily use cosine similarity as it is a widely adopted approach Sun et al. (2024); Zhang et al. (2021).
@@ 4.1 Weight Matrices of Experts
MoE models replace FFNs in standard transformers with experts. Following Geva et al. (2020); Qiu et al. (2024b), the projection matrices of the experts can be regarded as keys and values: the column vectors of $W_{\text{down}}$ represent potential outputs; the row vectors of $W_{\text{up}}$ produce weights for each possible output; the row vectors of $W_{\text{act}}$ determine whether to activate the corresponding neurons. Thus, examining the weight matrices provides a straightforward way to understand the expert behaviors. We analyze both the matrix and neuron levels to gain insights from different perspectives.
@@ 4.1.1 Matrix-level
In this part, we explore the similarity of the three projection matrices $W_{\text{up}}$, $W_{\text{act}}$, and $W_{\text{down}}$ among all experts in each layer. The similarity is calculated based on the flattened matrices and is illustrated in Fig. 1. We denote “F” as the Mistral FFN and “SE” as the DeepSeek shared expert. Note that the figures for different models do not share the same color bar.
@@
DeepSeek and Grok. The shared experts of DeepSeek are implemented as a single MLP block with a larger hidden size than the routed experts, preventing direct comparison of their flattened vectors; thus, we omit them from this experiment. Fig. 1 demonstrates that the similarities between the DeepSeek routed experts and Grok experts are close to zero. While Mixtrals’ training method remains unrevealed, it is known that DeepSeek and Grok are trained from scratch. This suggests that Mixtrals may have been trained using special schemes, resulting in less diverse experts compared to those trained from scratch Wu et al. (2022).
@@
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 1: Matrix-level similarity heat maps of expert weight matrices. Each layer contains three heat maps, corresponding to $W_{\text{up}}$, $W_{\text{act}}$, and $W_{\text{down}}$, respectively. The tick numbers refer to the expert indices. “F” denotes the Mistral FFN.
@@ 4.1.2 Neuron-level
In § 4.1, we measure the parameter similarity between experts at the matrix level. However, the calculation of cosine similarity is position-dependent. If the neurons of two experts are similar but in different orders, the similarity of their weight matrices will be significantly lower than expected. To address this, we propose two approaches to investigate the correlation at the neuron level: averaging and reordering. Averaging simply averages the rows (for $W_{\text{up}}$ and $W_{\text{act}}$) or the columns (for $W_{\text{down}}$) of the weight matrices, and then calculates the cosine similarity of the resulting vectors across experts. For reordering, we apply the Jonker-Volgenant algorithm Jonker and Volgenant (1988), which is typically used for solving linear assignment problems, to find the optimal order of neurons so that the cosine similarity between two experts is maximized.
We describe the results of the reordering method below and provide the details of the averaging approach in Append D. Additionally, the projection of neurons into low-dimensional spaces using PCA can be found in Append C.2. Due to the heavy computation, we only select several layers for the reordering calculation. Note that the matrices are reordered separately. We measure Kendall’s $\tau$ coefficient between the index sequences before and after reordering, whose value increases when the two sequences exist strong agreement. Tab. 2 depicts the common similarity growth after reordering and the average Kendall’s coefficient $\bar{\tau}$ over the selected layers. The order of Mixtral neurons changes little (resulting in a large $\tau$), and hence nearly unchanged similarities. Despite the substantial similarity increase for DeepSeek and Grok after reordering, their overall values remain around 1e-2.
@@ 4.2 Gate Embedding
The gate embedding of our investigated MoE models is implemented as a linear layer $W_{g}$ with size $\mathbb{R}^{N}\times\mathbb{R}^{d_{\text{hid}}}$, where $N$ is the number of experts. The gate serves as a crucial component of MoE, making it essential to study its attributes to understand MoE functionality better. In addition, since each row vector in the gate embedding determines expert selection, some correspondence may exist between $W_{g}$ and the expert weights.
To investigate this, we measure the similarities between the gate embedding vectors $W_{g}[n,:]$ for $n\in[1,N]$. For computational simplicity, we compare them with the neuron-level averaging (instead of the reordering) heat maps of experts presented in Append D, with qualitative analyses detailed in Append E. Specifically, we found that, for all four MoE models, the patterns in the heat maps of gate vectors and of expert neurons $W_{\text{act}}[i,:]$ are partially alike in some layers (i.e., the same coordinates in both heat maps exhibit relatively higher or lower values simultaneously).
@@
Model $(X,Y_{\text{up}})$ $(X,Y_{\text{act}})$ $(X,Y_{\text{down}})$
@@ 4.3 Summary
Here, we conclude the key observations from the analysis of static parameters: i) Mixtral might contain expert(s) with unique attributes, as evidenced by the frequent presence of dark crosses in Fig. 1. ii) The similarities of DeepSeek and Grok expert weight matrices are generally lower than those in Mixtrals. As mentioned in § 4.1.1, the matrix-level similarities of DeepSeek and Grok experts are typically close to zero, whereas Mixtrals’ expert similarities average around 0.3. iii) The weights of different experts become less similar in deeper layers, as observed in the Mixtrals’ heat maps in Fig. 1. iv) $W_{\text{up}}$, $W_{\text{down}}$, and $W_{\text{act}}$, share similar patterns in their similarity heat maps (Fig. 1). v) The similarities of $W_{g}$ and of $W_{\text{act}}$ show either positive or negative association. Tab. 3 depicts the $R_{\text{avg}}^{2}$ values, where the pairing of $W_{g}$ and $W_{\text{act}}$ achieves the highest correlation across all four models.
@@ 5 Analysis of Dynamic Behaviours
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption
Figure 2: Similarity heat maps of expert output features using the short input. The top $k$ experts for each token are shown on top of each heat map. The tick numbers refer to the expert indices. “F” and “SE” denote the Mistral FFN and the DeepSeek shared expert, respectively.
@@
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 3: Average similarity heat maps of expert output features using the long input, plotted along with the matrix-level similarity heat maps. The tick numbers refer to the expert indices. “F” denotes the Mistral FFN.
@@
Input data. We utilize a short input and a long input for the experiments in this section. For the short input, we employ the first few words of the input from another MoE-related work Cai et al. (2024)  [^4] The specific tokens are ¡s¿, As, an, open, source, alternative, to, where the start of the sentence symbol ¡s¿ does not applicable for the Grok tokenizer.. For the long input, we adopt 10 sequences from the test set of the WikiText-103 Merity et al. (2016) dataset, totaling approximately 1100 tokens. The sequences in WikiText-103 cover a variety of domains, with the 10 sequences we used spanning topics such as music, design, and construction. To ensure the robustness of our findings, we repeat experiments requiring the long input (§ 5.1, § 5.2) using additional datasets with over 80K tokens, including GSM8K Cobbe et al. (2021) and Magicoder- Evol-Instruct-110K Wei et al. (2024b). See Append F for details. The observations of these additional, subject-specific datasets align with the results described in the main context, demonstrating the universality of our conclusions.
@@ 5.1 Outputs of Experts
For clarity, the average similarity heat maps are plotted alongside the matrix-level similarity graphs of the expert weight matrices. Fig. 9 further depicts the results from additional datasets, which are consistent with those of the long input.
@@
DeepSeek. Given the short input, most similarities are around zero, while the values in the last layer are significantly larger. Again, the similarities between experts chosen by the gate are likely to be higher, although this difference occurs much less frequently than in Mixtrals. The average similarities for the long input also approach zero. Moreover, the number of “small rectangular” with relatively light color in the graphs decreases as the layer depth increases (except for the last layer), meaning that the average similarities gradually decline.
Grok. Surprisingly, the similarities between the output features remain high for all tokens in the short input, indicating the experts exhibit similar behaviours. However, the similarities of their weight matrices are mostly zeros (§ 4.1.1). We speculate that this may be due to the relatively large size of each Grok expert, allowing each to learn comprehensive knowledge and behave similarly despite having distinct parameters. When averaging the similarities for the long input, some of the resulting average heat maps display patterns similar to those of the $W_{\text{act}}$ figures. This relationship aligns with the observations made for Mixtrals.